# WaveCert: Formal Compiler Validation for Asynchronous Dataflow Programs

### v1

@@ Abstract.
Coarse-grained reconfigurable arrays (CGRAs), or dataflow architectures in general, have gained attention in recent years due to their promising power efficiency compared to traditional von Neumann architectures. To program these architectures using ordinary languages such as C, a dataflow compiler must transform the original sequential, imperative program into an equivalent dataflow graph, composed of dataflow operators running in parallel. This transformation is challenging since the asynchronous nature of dataflow graphs allows out-of-order execution of operators, leading to behaviors not present in the original imperative programs.
We address this challenge by developing a translation validation technique for dataflow compilers to ensure that the dataflow program has the same behavior as the original imperative program on all possible inputs and schedules of execution. We apply this method to a state-of-the-art dataflow compiler targeting the RipTide CGRA architecture. Our tool uncovers 7 compiler bugs where the compiler outputs incorrect dataflow graphs, including a data race that is otherwise hard to discover via testing. After repairing these bugs, our tool verifies the correct compilation of all programs in the RipTide benchmark suite.
@@ 1. Introduction
Even as Moore’s Law ends, many applications still demand better power efficiency, e.g., for high-performance machine learning or for ultra-low-power environmental monitoring. A flood of new hardware architectures have emerged to meet this demand, but one attractive approach that retains general-purpose programmability is a class of dataflow architectures called coarse-grained reconfigurable arrays (CGRAs) (cgra-survey, ). In a CGRA architecture, an array of processing elements (PEs) is physically laid out on the chip. Each PE can be configured to function as a high-level operator that performs an arithmetic operation, a memory operation, or a control-flow operation. These PEs are interconnected via an on-chip network, reducing communication and data movement costs compared to traditional von Neumann architectures, thus making a CGRA more power efficient.
Abstractly, CGRAs run dataflow programs, or networks of operators that execute independently and asynchronously. While CGRAs are able to run dataflow programs efficiently due to their inherent parallelism, this parallelism also introduces new, subtle correctness issues due to the potential for data races. As a result, dataflow programs are not written directly, but are instead compiled from ordinary, imperative languages. However, the gap between imperative and dataflow programming means that these compilers are complex; to maximize performance, they must use a variety of program analyses to convert imperative programs into equivalent distributed ones.
@@
In this work, we propose using translation validation (pnueli-tv, ) to certify the results of dataflow compilers. In our case, translation validation amounts to checking that the output dataflow program from the compiler has the same behaviors as the input imperative program. Compared to traditional testing, translation validation provides a much stronger guarantee, as it checks that the imperative and the dataflow programs behave the same on all possible inputs. Additionally, in contrast to full compiler verification, translation validation allows the use of unverified compilation toolchains, which is crucial for the fast-developing area of CGRA architectures.
@@
Our main challenge for leveraging translation validation is managing the asynchronicity of dataflow programs. Most translation validation works (pnueli-tv, ; necula-tv, ; kundu-tv, ; data-driven-eq, ) prove equivalence between input and output programs via (bi)simulation, or relational invariants between the states of the two programs. Since dataflow programs have exponentially more states than imperative programs (arising from asynchronous scheduling decisions), constructing this bisimulation relation directly would be quite tricky.
@@
Instead, we utilize a two-phase approach. First, we use symbolic execution to prove that there is a bisimulation between the input imperative program and the output dataflow program on a canonical schedule. In the canonical schedule, the dataflow operators are scheduled to fire in a similar order to their counterparts in the original imperative program.
@@
Next, we prove that any possible schedule of the dataflow program must converge to the same final state as the canonical schedule; i.e., it is confluent (term-rewriting, ). For our setting, proving confluence requires showing that the dataflow program does not contain any data races. Taking inspiration from separation logic (separation-logic, ), we augment our symbolic execution with linear permission tokens to track ownership of memory locations. Taken together with the initial bisimulation phase, we can prove that the imperative program is equivalent to the dataflow program on all possible inputs and schedules.
@@
We have implemented this technique in WaveCert  [^1] Available on GitHub at https://github.com/secure-foundations/riptide-verification., a translation validation system targeting the state-of-the-art CGRA architecture RipTide (riptide, ). RipTide features a dataflow compiler from LLVM (llvm, ) programs to dataflow programs. Using our tool, we found 7 compiler bugs where the RipTide dataflow compiler generated incorrect dataflow programs. All of these bugs were confirmed by the developer of the RipTide compiler. One of these bugs allows the compiler to emit a dataflow program with data races, which are hard to discover via testing and costly to fix after deployment.
@@
  • A novel two-phase translation validation technique for certifying compilation results from imperative programs to dataflow programs;
  • An implementation of our technique, WaveCert, which targets the RipTide dataflow compiler; and
  • An evaluation of WaveCert on the RipTide dataflow compiler, which revealed 7 compilation bugs. Each successful verification result takes less than 30 seconds, with most taking around 10 seconds.
@@
2. Preliminaries: CGRAs and RipTide
Dennis et al. (dennis-dataflow, ) introduced early designs for dataflow programming and architectures in the mid 1970s to allow programs to achieve a higher degree of concurrency. In the intervening years, the community has explored a wide variety of dataflow architectures (dataflow-architectures, ; yazdanpanah-dataflow, ). The desire for low-power, general-purpose programmability has recently led to a line of work on coarse-grained reconfigurable arrays (CGRAs) (cgra-survey, ).
@@ 2. Preliminaries: CGRAs and RipTide
To utilize the efficiency and parallelism in CGRAs, the predominant general-purpose way to program CGRAs is to compile from a sequential, imperative program into a dataflow graph via a dataflow compiler. This is used in CGRA projects such as PipeRench (piperench, ), WaveScalar (wavescalar, ), and more recently, RipTide (riptide, ) and Pipestitch (pipestitch, ).
In this work, we focus on the approach taken by RipTide (riptide, ), which has a dataflow compiler from arbitrary LLVM functions to dataflow programs. For an informal description of the semantics of the LLVM intermediate representation (IR), we refer the reader to the official LLVM IR manual (llvm-ir, ). In this work, we use the subset of LLVM IR that is supported by the RipTide compiler, including basic control-flow and integer arithmetic.
@@
The dataflow program output by the RipTide compiler is a Turing-complete representation of programs. It is represented as a dataflow graph in which the nodes are called operators and edges are called channels. Semantically, channels are unbounded FIFO queues used for communication between operators. Operators can be thought of as processes running in parallel with a local state. Repeatedly, each operator: waits to dequeue (a subset of) input channels to get input values; performs a local computation, optionally changing the local state or global memory; and pushes output values to (a subset of) output channels. When an iteration of this loop is done, we say that the operator has executed or fired.
We model the execution of dataflow operators using a fully asynchronous semantics. Different CGRAs would use different scheduling strategies (cgra-survey, ), so we conservatively use asynchronicity to cover all possible schedules.
@@
In Figure 1, we show the five most important types of operators. Reading from left to right, we first have the control flow operators, carry (C) and steer (T). The carry operator is used to support loop variables. It has two local states: in the initial state, it waits for a value from A, outputs it to O, and transitions to the loop state. In the loop state, it waits for values from B and D (decider); if D is true, it outputs O = B; otherwise it discards B and transitions to the initial state.
@@
The steer operator is used for conditional execution. It waits for values from A and D. If D is true, it outputs O = A; otherwise, it discards the value, and outputs nothing.
@@
To compile an LLVM program to a dataflow program, the RipTide compiler first maps each LLVM instruction to its corresponding dataflow operator, and then it enforces the expected control-flow and memory ordering semantics of the original program. Using a control dependency analysis, it inserts steer operators to selectively enable/disable operators depending on branch/loop conditions. Using a memory ordering analysis, it inserts dataflow dependencies between load/store operators to prevent data races.
These analyses are quite complex and use various optimizations to allow more parallelism than the original LLVM program. As shown in our evaluation in Section 6, this process can easily have bugs. Hence, in our work, we take the translation validation approach to certify the correctness of dataflow compilation.
Figure 2. An example of input/output programs from the RipTide dataflow compiler.
@@
Figure 2. An example of input/output programs from the RipTide dataflow compiler.
Figure 3. A trace of execution from the dataflow cut point 2 following the canonical schedule of operators: 1, 2, 4, 7, 8, 9, 6 (which are inferred from the LLVM trace from cut point 2 to cut point 2). In addition, steer operators 3 and 5 are also fired in step 3 immediately after they are ready. At the end of the trace, the configuration matches cut point 2 again. For each configuration, we show the state of the channels (solid lines for channels with one value, and dotted lines for empty channels). The lightning symbols indicate which operators are fired in the previous step. Before the firing of an operator, the input channels of the operator should be non-empty (i.e., marked with a solid line). After firing, the input channels are emptied, and the output channels will become non-empty. The exact symbolic values in each channel are omitted, but their attached permission variables (starting with p) are marked. At the top of each configuration, we also indicate the list of permission constraints we add at each step.
@@
Figure 3. A trace of execution from the dataflow cut point 2 following the canonical schedule of operators: 1, 2, 4, 7, 8, 9, 6 (which are inferred from the LLVM trace from cut point 2 to cut point 2). In addition, steer operators 3 and 5 are also fired in step 3 immediately after they are ready. At the end of the trace, the configuration matches cut point 2 again. For each configuration, we show the state of the channels (solid lines for channels with one value, and dotted lines for empty channels). The lightning symbols indicate which operators are fired in the previous step. Before the firing of an operator, the input channels of the operator should be non-empty (i.e., marked with a solid line). After firing, the input channels are emptied, and the output channels will become non-empty. The exact symbolic values in each channel are omitted, but their attached permission variables (starting with p) are marked. At the top of each configuration, we also indicate the list of permission constraints we add at each step.
@@
Index LLVM Dataflow Correspondence
1 %entry Initial config
$\mathrm{mem}_{\mathrm{llvm}}=\mathrm{mem}_{\mathrm{df}}$
2
%body
$\to$ %header
See Figure 3
$\mathrm{mem}_{\mathrm{llvm}}=\mathrm{mem}_{\mathrm{df}}$
@@
$\bot$ %end Final config
$\mathrm{mem}_{\mathrm{llvm}}=\mathrm{mem}_{\mathrm{df}}$
@@
3. Overview and an Example
WaveCert performs translation validation on the RipTide dataflow compiler, which compiles LLVM programs to dataflow programs. Given the input program and the compiled output program (along with some hints generated by the compiler), WaveCert performs two checks: first, a bisimulation check, which verifies that the LLVM program is equivalent to the dataflow program on a restricted, canonical schedule of dataflow operators; and second, a confluence check, which proves that the choice of schedule for the dataflow program does not affect the final result.
If a pair of input/output programs passes both checks, the two programs are equivalent. If the bisimulation check fails, there might be a compilation bug that produced an incorrect dataflow. If the bisimulation check succeeds but the confluence check fails, there might be a data race in the output dataflow program.
Example
In Figure 1(b), we have an example input and output program from the dataflow compiler. The input LLVM function @test in Figure 1(a) contains a single loop with loop header block %header and loop body block %body. The loop increments variable %i from 0 to %len, and at each iteration, updates %B[%i] = %A[%i] + 1.
The LLVM function @test gets compiled to the dataflow program in Figure 1(b), in which most operators correspond to an LLVM instruction:
  • The comparison operator 4 corresponds to line 7;
  • The load operator 7 corresponds to line 12;
  • The add operator 8 corresponds to line 13;
  • The store operator 9 corresponds to line 15; and
  • The add operator 6 corresponds to line 16.
The compiler also inserts additional operators to faithfully implement the sequential semantics of the LLVM program. The steer operators marked with T enforce that the operators corresponding to the loop body only execute when the loop condition is true. The carry operator 1 is used for the loop induction variable %i (line 6). The carry operator 2 is used for a loop variable inserted by the compiler to enforce a memory dependency: since arrays A and B could overlap, we have an additional dataflow path from the store operator, passing through the carry operator 2 and the steer operator 5, to the load operator, in order to enforce that the load in the next iteration waits until the previous store has finished.
Bisimulation Check
The first step in WaveCert’s translation validation procedure is to check that the LLVM function is bisimilar to the dataflow program on a canonical schedule of operators, where the operators are executed in the same order as their LLVM instruction counterparts.
To check a bisimulation, we need to construct a relation between the states of the LLVM and dataflow programs. WaveCert constructs this relation by placing corresponding cut points on both the LLVM and the dataflow sides. A cut point symbolically describes a set of pairs of LLVM and dataflow configurations satisfying a particular correspondence condition, and a list of cut points together describes the bisimulation relation. Figure 4 shows the list of cut points for the example. WaveCert places cut point 1 for the initial configurations, cut point 2 for the loop structure, and a final cut point $\bot$ for all final configurations. For each pair of LLVM and dataflow cut points, WaveCert also infers a list of equations expressing the correspondence of symbolic variables in these cut points.
To automatically check that the proposed relation in Figure 4 is indeed a bisimulation, we perform symbolic execution from all pairs of cut points (except for $\bot$) until they reach another pair of cut points. Symbolic execution (symbolic-execution, ) is similar to concrete execution of programs, except with concrete values replaced by symbolic expressions. In this way, we can execute all possible concrete instances of a cut point altogether. On the LLVM side, we first symbolically execute cut point 1, which turns into two symbolic branches: reaching %header after one loop iteration (cut point 2), or failing the loop condition and reaching %end (cut point $\bot$). For LLVM cut point 2, we also have two branches: one reaching cut point 2 again after a loop iteration, the other failing the loop condition and reaching cut point $\bot$.
For the dataflow side, we have a correspondence between a subset of LLVM instructions and a subset of dataflow operators, which is automatically generated by the dataflow compiler. We use this mapping to infer the order in which we fire the dataflow operators, i.e., the canonical schedule. For example, from LLVM cut point 2, there are two branches going to cut point 2 and cut point $\bot$. These two branches have two traces of LLVM instructions (excluding br):
  • For the first branch to cut point 2: lines 6, 7, 11 - 16;
  • For the second branch to cut point $\bot$: lines 6, 7.
For the first branch, the list of LLVM instructions map to dataflow operators 1, 2, 4, 7, 8, 9, 6 (excluding ones that do not have a corresponding operator; operator 2 corresponds to a PHI instruction implicitly inserted during compilation for memory ordering). LLVM instructions in the second branch map to dataflow operators 1, 2, and 4. Therefore, for dataflow cut point 2, we execute these two traces of dataflow operators (and also any steer operator that can be executed), which gives us two dataflow configurations. Figure 3 shows the first branch from cut point 2 to itself. At the end of the execution, we can see that the configuration matches the cut point 2 configuration we started with, except with potentially different memory configurations and symbolic variables (e.g. i + 1 vs. i in the channel from operator 6 to 1).
@@ 3. Overview and an Example
$\mathrm{mem}_{\mathrm{llvm}}[\texttt{\%B}[\texttt{\%i}]\mapsto\texttt{\%A}[\texttt{\%i}]+1]=\mathrm{mem}_{\mathrm{df}}[\texttt{B}[\texttt{i}]\mapsto\texttt{A}[\texttt{i}]+1]$
%i + 1=i + 1%i + $=\texttt{i + 1}$ i + 1
If this check succeeds for all dataflow and LLVM cut points, we soundly conclude that we have obtained a bisimulation between the LLVM program and the dataflow program (under the deterministic canonical schedule).
@@
Once we have established a bisimulation between the LLVM function and the canonical schedule of the dataflow program, we check that the dataflow program is confluent: all possible schedules converge to the canonical schedule in the final state (or synchronize with it infinitely often if the program is non-terminating).
@@
We attach a permission variable (representing a permission token that we do not know a priori) to each value in the channels of dataflow configurations in each symbolic branch. Figure 3 shows an example of this attachment of permission variables and permission constraints. At the starting state cut point 2, we attach free permission variables $p_{1},p_{2},p_{3},p_{4}$ to values in the configuration. After the first step in which operators 1 and 2 fire, these permission variables are consumed by the two carry operators and we generate fresh permission variables $p_{5},p_{6},p_{7}$ for the new output values. We require that the permission variables are used in a linear fashion: operators cannot duplicate or generate permission tokens, instead, they have to be obtained from the inputs. So for the first step in Figure 3, we add two constraints $p_{5}+p_{6}\leq p_{1}+p_{3}$ and $p_{7}\leq p_{2}+p_{4}$ to say that the output permissions have be contained in the input permissions. Furthermore, if the operator is a load or a store, such as in Figure 3 steps 4 and 6, we require that the suitable permission (read for load, and write for store) should be present in the input permissions (e.g. $\texttt{write}\leq p_{13}+p_{17}$ at step 6 in Figure 3).
@@
Finally, when the execution is finished and has either hit a final state or another cut point, we have collected a list of constraints involving a set of permission variables. If the configuration is matched to a cut point, we add additional constraints to make sure that the assignment of tokens at the cut point is consistent. For example in Figure 3, for the last configuration, we would add these additional constraints: $p_{1}=p_{19}$, $p_{2}=p_{18}$, $p_{3}=p_{8}$, $p_{4}=p_{9}$. We then solve for the satisfiability of these constraints with respect to the permission algebra ($(\{0,\texttt{read},\texttt{write}\},+,\leq)$ we have defined above.

### v2

# WaveCert: Translation Validation for Asynchronous Dataflow Programs via Dynamic Fractional Permissions

@@ Abstract.
Coarse-grained reconfigurable arrays (CGRAs) have gained attention in recent years due to their promising power efficiency compared to traditional von Neumann architectures. To program these architectures using ordinary languages such as C, a dataflow compiler must transform the original sequential, imperative program into an equivalent dataflow graph, composed of dataflow operators running in parallel. This transformation is challenging since the asynchronous nature of dataflow graphs allows out-of-order execution of operators, leading to behaviors not present in the original imperative programs.
We address this challenge by developing a translation validation technique for dataflow compilers to ensure that the dataflow program has the same behavior as the original imperative program on all possible inputs and schedules of execution. We apply this method to a state-of-the-art dataflow compiler targeting the RipTide CGRA architecture. Our tool uncovers 8 compiler bugs where the compiler outputs incorrect dataflow graphs, including a data race that is otherwise hard to discover via testing. After repairing these bugs, our tool verifies the correct compilation of all programs in the RipTide benchmark suite.
copyright: none
@@ 1. Introduction
Even as Moore’s Law ends, many applications still demand better power efficiency, e.g., for high-performance machine learning or for ultra-low-power environmental monitoring. A flood of new hardware architectures has emerged to meet this demand, but one attractive approach that retains general-purpose programmability is a class of dataflow architectures called coarse-grained reconfigurable arrays (CGRAs) (Liu et al., 2019). In a CGRA architecture, an array of processing elements (PEs) is physically laid out on the chip. Each PE can be configured to function as a high-level operator that performs an arithmetic operation, a memory operation, or a control-flow operation. These PEs communicate via an on-chip network, reducing data movement costs compared to traditional von Neumann architectures, thus making a CGRA more power efficient.
Abstractly, CGRAs run dataflow programs, or networks of operators that execute independently and asynchronously. While CGRAs are able to run dataflow programs efficiently due to their inherent parallelism, this parallelism also introduces new, subtle correctness issues due to the potential for data races. As a result, dataflow programs are not written directly, but are instead compiled from sequential, imperative languages. However, the gap between imperative and dataflow programming means that these compilers are complex; to maximize performance, they must use a variety of program analyses to convert imperative programs into equivalent distributed ones.
@@
In this work, we propose using translation validation (Pnueli et al., 1998) to certify the results of dataflow compilers. In our case, translation validation amounts to checking that the output dataflow program from the compiler has the same behaviors as the input imperative program. Compared to traditional testing, translation validation provides a much stronger guarantee, as it checks that the imperative and the dataflow programs behave the same on all possible inputs. Additionally, in contrast to full compiler verification, translation validation allows the use of unverified compilation toolchains, which is crucial for the fast-developing area of CGRA architectures.
@@
Our main challenge for leveraging translation validation is managing the asynchrony of dataflow programs. Most translation validation works (Pnueli et al., 1998; Necula, 2000; Kundu et al., 2009; Sharma et al., 2013) prove equivalence between input and output programs via simulation, or relational invariants between the states of the two programs. Since dataflow programs have exponentially more states than imperative programs (arising from asynchronous scheduling decisions), constructing this simulation relation directly would be quite tricky.
@@
Instead, we utilize a two-phase approach. First, we use symbolic execution to prove that there is a simulation between the input imperative program and the output dataflow program on a canonical schedule. In the canonical schedule, the dataflow operators are scheduled to fire in a similar order to their counterparts in the original imperative program.
@@
Next, we prove that any possible schedule of the dataflow program must converge to the same final state as the canonical schedule; i.e., it is confluent (Baader and Nipkow, 1998). For our setting, proving confluence requires showing that the dataflow program does not contain any data races. Inspired by fractional permissions (Boyland, 2013), we augment our symbolic execution with linear permission tokens to track ownership of memory locations. In contrast with interactive program logics (e.g., concurrent separation logic (Brookes, 2006)) that typically require manual annotations, we automatically compute dynamic permissions that may flow in arbitrary (sound) ways.
Putting together both phases, we prove that the imperative program is equivalent to the dataflow program on all possible inputs and schedules. Furthermore, equivalence with the sequential program also implies liveness and deadlock freedom for the dataflow program, as it enforces that the dataflow program can make progress whenever the sequential program can.
@@
We have implemented this technique in WaveCert, a translation validation system targeting the state-of-the-art CGRA architecture RipTide (Gobieski et al., 2023), which operates via a compiler from LLVM (Lattner and Adve, 2004) programs to dataflow programs. Using our tool, we found 8 compiler bugs where the RipTide dataflow compiler generated incorrect dataflow programs. All of these bugs were confirmed by the developers of the RipTide compiler. One of these bugs allows the compiler to emit a dataflow program with data races, which are hard to discover via testing and costly to fix after deployment.
@@
  • A novel two-phase translation validation technique to prove that the output dataflow program has equivalent behavior as the input imperative program, capturing both correctness and liveness properties.
  • An implementation of our technique, WaveCert  [^1] Source code is available in an GitHub repository (Lin et al., 2024), targeting the RipTide dataflow compiler.
  • An evaluation of WaveCert on the RipTide dataflow compiler, which reveals 8 compilation bugs. Most verification result takes around 10 seconds with a maximum of about 30 seconds.
[…]
@@
2. Preliminaries: Dataflow Programs, CGRAs, and RipTide
Dataflow programming (Johnston et al., 2004) is an old idea dating back to the 1960s and 1970s. Early foundational works include the dataflow architecture by Dennis et al. (Dennis and Misunas, 1974), as well as theoretical analysis of models of dataflow programs such as computation graphs (Karp and Miller, 1966) and Kahn process networks (Gilles, 1974). The main motivation in these early works is to achieve a greater degree of instruction-level parallelism (Ackerman, 1982).
In the intervening years, the community has explored a wide variety of dataflow architectures (Arvind and Culler, 1986; Yazdanpanah et al., 2014). Among these dataflow architectures, coarse-grained reconfigurable arrays (CGRAs) (Liu et al., 2019) stand out due to their promising power efficiency: recent work (Gobieski et al., 2021, 2023) is able to support general programmability, while reducing the energy cost by orders of magnitude compared to a von Neumann core.
@@ 2. Preliminaries: Dataflow Programs, CGRAs, and RipTide
To utilize the efficiency and parallelism in CGRAs, the predominant general-purpose way to program CGRAs is to compile a sequential, imperative program into a dataflow graph via a dataflow compiler, as seen in CGRA projects such as PipeRench (Goldstein et al., 2000), WaveScalar (Swanson et al., 2003), and more recently, RipTide (Gobieski et al., 2023) and Pipestitch (Serafin et al., 2023).
In this work, we focus on the approach taken by RipTide (Gobieski et al., 2023), which has a dataflow compiler from arbitrary LLVM functions to dataflow programs. We briefly introduce LLVM and dataflow programs below.
LLVM IR
LLVM (Lattner and Adve, 2004) is a compiler framework and features an intermediate representation called LLVM IR. Compiler writers can translate a higher-level language such as C or Rust into LLVM IR, and then the LLVM framework can perform optimizations on LLVM IR and produce machine code in various target architectures. For an informal description of the semantics of the LLVM IR, we refer the reader to the official LLVM IR manual (LLVM, 2024b). In this work, we use the subset of LLVM IR supported by the RipTide compiler, including basic control-flow, integer arithmetic, and memory operations. Function calls and floating-point operations are not supported.
Dataflow programs
@@
A dataflow program is a Turing-complete representation of programs. It is represented as a dataflow graph in which the nodes are called operators and edges are called channels. Semantically, operators can be thought of as stateful processes running in parallel, which communicate through channels, or FIFO queues of messages. Repeatedly, each operator: waits to dequeue (a subset of) input channels to get input values; performs a local computation, optionally changing its local state or global memory; and pushes output values to (a subset of) output channels. When an iteration of this loop is done, we say that the operator has executed or fired. Since different CGRAs have different strategies for scheduling operators (Liu et al., 2019), we conservatively use asynchrony to model all possible schedules.
@@
In Figure 1, we show the five most important types of operators in RipTide (Gobieski et al., 2023) [^2] WaveCert supports all RipTide operators, except for the Stream operator, as described in Section 7.. In Figure 1, reading from left to right, we first have the control flow operators, carry (C) and steer (T). The carry operator is used to support loop variables. It has two local states: in the initial state, it waits for a value from A, outputs it to O, and transitions to the loop state. In the loop state, it waits for values from B and D (decider); if D is true, it outputs O = B; otherwise it discards B and transitions to the initial state.
@@
The steer operator is used for conditional execution. It waits for values from A and D. If D is true, it outputs O = A; otherwise, it discards the value and outputs nothing.
@@
To compile an LLVM program to a dataflow program, the RipTide compiler first maps each LLVM instruction to its corresponding dataflow operator, and then it enforces the expected control-flow and memory ordering semantics of the original program. Using a control dependency analysis, it inserts steer operators to selectively enable/disable operators depending on branch/loop conditions; Using a memory ordering analysis, it inserts dataflow dependencies between load/store operators to prevent data races.
These analyses are quite complex and use various optimizations to allow more parallelism than the original LLVM program. As shown in our evaluation in Section 6, this process can easily have bugs. Hence, in our work, we use translation validation to certify the correctness of dataflow compilation.
3. Overview and an Example
WaveCert performs translation validation on the RipTide dataflow compiler, which compiles LLVM programs to dataflow programs. Given the input program and the compiled output program (along with some hints generated by the compiler), WaveCert performs two checks: first, a simulation check, which verifies that the LLVM program is equivalent to the dataflow program on a restricted, canonical schedule of dataflow operators; and second, a confluence check, which proves that the choice of the canonical schedule for the dataflow program is general and all other schedules reach the same final state as the canonical schedule.
@@ 3. Overview and an Example
Figure 2. An example of input/output programs from the RipTide dataflow compiler. Most dataflow operators correspond to LLVM instructions: the comparison operator corresponds to line 6, the add operators correspond to lines 11 and 14, while the load and store operators correspond to lines 10 and 13, respectively.
[…]
@@
If a pair of input/output programs passes both checks, the two programs are equivalent (in the sense of Definition 2). Equivalence implies both safety — the dataflow program returns the correct values — and liveness — the dataflow program always terminates correctly, and does not contain communication-related deadlocks.
Example
In Figure 2(b), we have an example input and output program from the dataflow compiler. The input LLVM function @test in Figure 2(a) contains a single loop with loop header block %header and loop body block %body. The loop increments variable %i from 0 to %len, and at each iteration, updates %B[%i] = %A[%i] + 1. The LLVM function @test is compiled to the dataflow program in Figure 2(b).
In addition to compute and memory operators, the compiler also inserts operators to faithfully implement the sequential semantics of the LLVM program. The steer operators marked with T enforce that the operators corresponding to the loop body only execute when the loop condition is true. The carry operator 1 is used for the loop induction variable %i (line 5). The carry operator 2 is used for a loop variable inserted by the compiler to enforce a memory dependency: since arrays A and B could overlap, we have an additional dataflow path along operators $9\to 2\to 5\to 7$ in order to enforce that the load in the next iteration waits until the previous store has finished.
Figure 3. This plot shows a trace of execution from dataflow cut point 2 (upper left corner). The operators fired in this trace are 1, 2, 4, 3, 5, 7, 8, 9, 6 (which are inferred from the LLVM trace from cut point 2 to cut point 2). At the end of the trace, the configuration matches cut point 2 again. In each configuration, we mark the fired operators with orange lightning symbols. The emptiness of each channel after firing the operator(s) is represented by dashed vs solid lines. Before the firing of an operator, the input channels of the operator should be non-empty and marked with a solid line. After firing, the input channels are emptied, and the modified output channels will become non-empty and marked with a solid, orange line. The symbolic values in each channel are omitted, but their attached permission variables (starting with p) are marked. At the top of each configuration, we also indicate the permission constraints added at each step.
[…]
@@
Index LLVM Program Point Dataflow Configuration Correspondence
1 %entry Initial config $\mathsf{Mem}_{\mathrm{llvm}}=\mathsf{Mem}_{\mathrm{df}}$
@@
$\bot$ %end Final config $\mathsf{Mem}_{\mathrm{llvm}}=\mathsf{Mem}_{\mathrm{df}}$
@@
[…]
@@
Simulation Check
The first step in WaveCert’s translation validation procedure is to check that the LLVM and dataflow program are equivalent on a canonical schedule of dataflow operators, where the dataflow operators are executed in the same order as their LLVM instruction counterparts.
This is done using a simulation relation, or relational invariant, between the states of LLVM and dataflow programs. WaveCert constructs this relation by placing corresponding cut points on both the LLVM and the dataflow sides. A cut point symbolically describes a set of pairs of LLVM and dataflow configurations satisfying a correspondence condition; a list of cut points together describes the simulation relation. Figure 4 shows the list of cut points for the example. WaveCert places cut point 1 for the initial configurations, cut point 2 for the loop structure, and a final cut point $\bot$ for all final configurations. For each pair of LLVM and dataflow cut points, WaveCert also infers a list of equations expressing the correspondence of symbolic variables in these cut points.
To automatically check that the proposed relation in Figure 4 is indeed a simulation, we perform symbolic execution (Baldoni et al., 2018) from all pairs of cut points (except for $\bot$) until they reach another pair of cut points. On the LLVM side, we first symbolically execute cut point 1, which turns into two symbolic branches: reaching %header after one loop iteration (cut point 2), or failing the loop condition and reaching %end (cut point $\bot$). For LLVM cut point 2, we also have two branches: one reaching cut point 2 again after a loop iteration, the other failing the loop condition and reaching cut point $\bot$.
For the dataflow side, we have a correspondence between a subset of LLVM instructions and a subset of dataflow operators, which is automatically generated by the dataflow compiler. We use this mapping to infer the order in which we fire the dataflow operators, i.e., the canonical schedule. For example, from LLVM cut point 2, there are two branches reaching cut point 2 and cut point $\bot$ respectively. These two branches have two traces of LLVM instructions (excluding br):
  • For the first branch to cut point 2: lines 5, 6, 9 - 14;
  • For the second branch to cut point $\bot$: lines 5, 6.
For the first branch, the LLVM instructions map to dataflow operators 1, 2, 4, 7, 8, 9, 6 (excluding ones that do not have a corresponding operator; operator 2 corresponds to a phi instruction implicitly inserted during compilation for memory ordering). LLVM instructions in the second branch map to dataflow operators 1, 2, and 4. Therefore, for dataflow cut point 2, we execute these two traces of dataflow operators (and also any steer operator that can be executed), which gives us two dataflow configurations. Figure 3 shows the trace of execution of the first branch from cut point 2 to itself. At the end of the execution, we can see that the configuration matches the cut point 2 we started with, except with a different memory state and symbolic variables (e.g. i + 1 vs. i in the channel from operator 6 to 1).
@@
If this check succeeds for all cut points, we soundly conclude that we have obtained a simulation relation between the LLVM program and the dataflow program on the canonical schedule.
@@
Once we have established a simulation relation between the LLVM program and a canonical schedule of the dataflow program, we check that the dataflow program is confluent: all possible schedules converge to the canonical schedule in the final state (or synchronize with it infinitely often if the program is non-terminating). This confluence result ensures that our simulation result for the canonical schedule is fully general for all other schedules.
@@
In general, as defined in Section 5, we allow write to split into $k$ tokens of read, for a predetermined value of $k$, and we have one permission token for every memory region, such as the array A or B. The intuition is that to write to a memory location, an operator needs to have exclusive write ownership of that location and no other operator should have write or read permissions; while to read a memory location, we allow potentially $k$ parallel reads to happen independently. If all reads are done, and a write needs to occur, we merge all read tokens into one write token.
These permission tokens are passed between operators via channels, and operators are not allowed to duplicate or create new tokens (i.e., permission tokens are used linearly). We attach permissions to values flowing through the dataflow program, instead of creating separate channels for communicating permission tokens.
@@
To determine the exact assignment and flow of permission tokens, we first attach a permission variable (representing a permission token that we do not know a priori) to each value in the channels of dataflow configurations in each symbolic branch. Figure 3 shows an example of this attachment of permission variables and permission constraints. At the starting state cut point 2, we attach free permission variables $p_{1},p_{2},p_{3},p_{4}$ to values in the configuration. After the first step in which operators 1 and 2 fire, these permission variables are consumed by the two carry operators and we generate fresh permission variables $p_{5},p_{6},p_{7}$ for the new output values. We require that the permission variables are used in a linear fashion: operators cannot duplicate or generate permission tokens; instead, they have to be obtained from the inputs. So for the first step in Figure 3, we add two constraints $p_{5}+p_{6}\leq p_{1}+p_{3}$ and $p_{7}\leq p_{2}+p_{4}$ to say that the output permissions have be contained in the input permissions. Furthermore, if the operator is a load or a store, such as in Figure 3 steps 4 and 6, we require that the suitable permission (read for load, and write for store) is present in the input permissions (e.g. $\texttt{write}\leq p_{13}+p_{17}$ at step 6 in Figure 3).
@@
Finally, when the execution is finished and has either hit a final state or another cut point, we have collected a list of constraints involving a set of permission variables. If the configuration is matched to a cut point, we add additional constraints to make sure that the assignment of tokens at the cut point is consistent. For example in Figure 3, since the last configuration is matched again to the same cut point 2, we would add these additional constraints: $p_{1}=p_{19}$, $p_{2}=p_{18}$, $p_{3}=p_{8}$, $p_{4}=p_{9}$. We then solve for the satisfiability of these constraints with respect to the permission algebra ($(\{0,\texttt{read}_{1},\texttt{read}_{2},\texttt{write}\},+,\leq)$ that we have defined above.